If the two legs of a right triangle are 5m and 7m respectively, what is the hypotenuse? Round answer to two decimals.


1
Expert's answer
2021-02-06T10:05:13-0500

Let "a = 5m, b = 7m", and "c" is the hypotenuse. By Pythagorean's theorem, the hypotenuse is:


"c = \sqrt{a^2 + b^2}\\
c = \sqrt{5^2 + 7^2} = \sqrt{74}m\approx 8.6m"

Answer. "\sqrt{74}m" or "8.6m".